Differentiate from first principle square root of (x+√x).
1
Expert's answer
2019-10-21T10:28:34-0400
"f(x)=\sqrt{x+\sqrt{x}}"


"f'(x)=\bigg(\sqrt{x+\sqrt{x}}\bigg)'=\lim\limits_{h\rarr0}{f(x+h)-f(x) \over h}="

"=\lim\limits_{h\rarr0}{\sqrt{x+h+\sqrt{x+h}}-\sqrt{x+\sqrt{x}} \over h}="

"=\lim\limits_{h\rarr0}{\sqrt{x+h+\sqrt{x+h}}-\sqrt{x+\sqrt{x}} \over h}\cdot{\sqrt{x+h+\sqrt{x+h}}+\sqrt{x+\sqrt{x}} \over \sqrt{x+h+\sqrt{x+h}}}="

"=\lim\limits_{h\rarr0}{ x+h+\sqrt{x+h}-x-\sqrt{x}\over h(\sqrt{x+h+\sqrt{x+h}}+\sqrt{x+\sqrt{x}})}="


"=\lim\limits_{h\rarr0}{ h\over h(\sqrt{x+h+\sqrt{x+h}}+\sqrt{x+\sqrt{x}})}+"


"+\lim\limits_{h\rarr0}{ \sqrt{x+h}-\sqrt{x}\over h(\sqrt{x+h+\sqrt{x+h}}+\sqrt{x+\sqrt{x}})}="

"={ 1\over \sqrt{x+0+\sqrt{x+0}}+\sqrt{x+\sqrt{x}}}+"


"+\lim\limits_{h\rarr0}{ \sqrt{x+h}-\sqrt{x}\over h(\sqrt{x+h+\sqrt{x+h}}+\sqrt{x+\sqrt{x}})}\cdot{\sqrt{x+h}+\sqrt{x} \over \sqrt{x+h}+\sqrt{x}}="

"={ 1\over 2\sqrt{x+\sqrt{x}}}+\lim\limits_{h\rarr0}{ x+h-x\over h(\sqrt{x+h+\sqrt{x+h}}+\sqrt{x+\sqrt{x}})(\sqrt{x+h}+\sqrt{x})}="


"={ 1\over 2\sqrt{x+\sqrt{x}}}+\lim\limits_{h\rarr0}{ 1\over (\sqrt{x+h+\sqrt{x+h}}+\sqrt{x+\sqrt{x}})(\sqrt{x+h}+\sqrt{x})}="

"={ 1\over 2\sqrt{x+\sqrt{x}}}+{ 1\over (\sqrt{x+0+\sqrt{x+0}}+\sqrt{x+\sqrt{x}})(\sqrt{x+0}+\sqrt{x})}="

"={ 1\over 2\sqrt{x+\sqrt{x}}}\big(1+{1 \over 2\sqrt{x}}\big)"



"\bigg(\sqrt{x+\sqrt{x}}\bigg)'={ 1\over 2\sqrt{x+\sqrt{x}}}\bigg(1+{1 \over 2\sqrt{x}}\bigg)"